# HiFi-Mamba: Dual-Stream $\mathcal{W}$-Laplacian Enhanced Mamba for High-Fidelity MRI Reconstruction

### v1

@@ Abstract
I Introduction
Magnetic Resonance Imaging (MRI) is a clinically indispensable modality due to its non-invasive nature and excellent soft-tissue contrast [1, 2]. However, a primary limitation lies in its long acquisition time, which can cause patient discomfort and increase the risk of motion artifacts [3, 4]. To accelerate scans, modern protocols commonly adopt undersampling in the frequency domain. While this reduces scan time, it violates the Nyquist-Shannon sampling criterion [5, 6], leading to aliasing artifacts and degraded image quality. Addressing this challenge requires advanced reconstruction methods capable of recovering high-fidelity images from incomplete k-space data. Traditional compressed sensing (CS) methods [7] exploit sparsity priors in transform domains (e.g., wavelets), but they require extensive hyperparameter tuning and often lack robustness to variations in sampling patterns or noise conditions.
@@ I Introduction
Recent deep learning frameworks have substantially advanced MRI reconstruction by leveraging data-driven priors from large-scale datasets [8]. Convolutional neural networks (CNNs) have demonstrated strong performance by modeling hierarchical and localized anatomical features [9, 10]. Model-based CNNs further improve reconstruction quality by integrating the MRI forward model and enforcing data consistency [3, 11, 12]. However, the inherent locality of CNNs limits their capacity to capture long-range dependencies, which are crucial for reconstructing global anatomical structures, especially under highly undersampled conditions.
@@
Transformer-based architectures model global dependencies through self-attention by computing pairwise correlations across all spatial tokens [13]. This capacity has shown promise in MRI reconstruction [14, 15] by enabling global context modeling to restore anatomical structures. Nonetheless, standard Transformers incur quadratic complexity, posing challenges for high-resolution MRI applications. To improve efficiency, variants such as the Swin Transformer [16] employ hierarchical designs and shifted window-based attention to restrict computations to local neighborhoods. While computationally efficient, these approaches inherently constrain the receptive field, reigniting the fundamental trade-off between computational efficiency and global modeling capability, leaving a critical gap for a more holistic solution.
@@
Structured State Space Models (SSMs) [17] have recently emerged as promising alternatives to self-attention for global context modeling, offering linear-time scalability with strong sequence modeling capacity. Among them, Mamba [18] employs input-dependent state transitions and efficient gating mechanisms, enabling expressive long-range interactions with reduced computational complexity. While initially proposed for language tasks, Mamba is rapidly emerging as a powerful alternative in visual domains such as image classification and restoration [19, 20].
@@
However, a direct application of Mamba to MRI reconstruction reveals three fundamental limitations rooted in its original design. First, existing vision-specific Mamba architectures utilize multi-directional spatial scanning [21, 22] to enhance coverage, but this introduces significant computational redundancy (see Figure 1(a)). Second, its state-space parameters are derived independently for each spatial token through local transformations, limiting spatial awareness. This design neglects neighboring context, which is essential for modeling coherent anatomical structures in MRI. Third, due to its continuous-time formulation and discretization process that naturally favors smoother signal representations, existing Mamba-based variants lack sensitivity to high-frequency components [23], which are critical for preserving fine anatomical details in MRI reconstruction.
@@ II Related Work
CNNs have been widely employed in MRI reconstruction for their ability to extract hierarchical features efficiently. Early models such as DeepCascade [9] and ISTA-Net [24] framed reconstruction as an unrolled optimization process, integrating data fidelity with learnable, network-based priors. Subsequent approaches, including KIKI-Net [25] and DuDoRNet [26], incorporated architectural advances such as residual connections, dilated convolutions, and hybrid modeling across image and k-space domains to enhance reconstruction quality and optimization stability. However, the inherently localized receptive fields of CNNs constrain their ability to capture global anatomical context. This limitation has motivated the development of architectures with enhanced global modeling capacity.
@@
Transformers have been increasingly adopted in MRI reconstruction for their ability to model long-range dependencies through global self-attention, which facilitates the preservation of anatomical structures. Early models such as SLATER [27] and TTM [28] applied Transformer blocks in the image domain. DuDReTLU-Net [29] jointly models image and $k$-space domains through a transformer-based architecture to enhance reconstruction quality. Later variants such as SwinMR [30] and ReconFormer [14] adopted hierarchical and windowed self-attention mechanisms to improve multi-scale feature modeling and computational efficiency. Although these designs reduce computational overhead, they often rely on localized attention and staged aggregation, which may still limit the capacity for global context modeling in high-resolution MRI reconstruction.
Structured State Space Models and Mamba
State space models (SSMs) have recently gained attention for their ability to model long-range dependencies with linear complexity [31]. The vision-centric Mamba variant, vMamba [20], introduced a four-directional scanning strategy to enhance 2D spatial context modeling. Recent works, such as MambaMIR [32] and LMO [33], have extended this paradigm for MRI reconstruction. However, these adaptations remain suboptimal in both efficiency and fidelity. Their reliance on multi-directional scanning induces considerable computational overhead due to repeated processing of spatial features—an issue exacerbated in high-resolution scenarios. Moreover, they inherit two core limitations of the original Mamba design: purely local state-space parameterization and insensitivity to high-frequency anatomical details. These constraints highlight a gap between current vision-based Mamba variants and the spectral characteristics of MRI.
III Methodology
III-A Overall Pipeline
Our proposed HiFi-Mamba network follows an unrolled optimization framework, a powerful paradigm for solving inverse problems like MRI reconstruction. The network backbone consists of $K=6$ cascaded HiFi-Mamba Groups. As shown in Figure 2, each Group consists of two sequential Mamba Units for feature refinement, followed by a Data Consistency (DC) block. Specifically, the input undersampled image $I_{\text{in}}\in\mathbb{R}^{H\times W\times 2}$ is a two-channel tensor representing the real and imaginary parts of the complex-valued MR data. The input tensor is first transformed into patch embeddings as:
@@ III-B The Mamba Unit
III-C $\mathcal{W}$-Laplacian Block
@@
While the Laplacian pyramid [34] offers residual-based multiscale representations, its decomposition is resolution-oriented rather than frequency-structured. The resulting low-frequency component is a blurred, downsampled approximation without explicit spectral semantics, limiting its utility for preserving anatomical context in MRI. To overcome this limitation, we replace the resolution-based hierarchy with a wavelet-based formulation that enables structured and reversible frequency separation.
@@
III-D HiFi-Mamba Block
@@
Figure 3: Qualitative comparison on the fastMRI and CC359 datasets under single-coil settings. (a) Reconstruction results on the fastMRI knee dataset with acceleration factors AF=4 and AF=8. (b) Reconstruction results on the CC359 brain dataset under the same acceleration factors. The second row of each subplot shows the corresponding error maps. The blue boxes, yellow ellipses and red arrow highlight the details in the reconstruction results.
@@
TABLE I: Quantitative comparison on the fastMRI-Equispaced and CC359-Equispaced under $4\times$ and $8\times$ acceleration factors.
@@
PSNR $\uparrow$ SSIM $\uparrow$ NMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ NMSE $\downarrow$
AF=4 AF=8 AF=4 AF=8 AF=4 AF=8 AF=4 AF=8 AF=4 AF=8 AF=4 AF=8
@@
UNet [8] 31.66 28.60 0.798 0.697 0.021 0.035 28.27 24.28 0.847 0.720 0.025 0.059
Ista [24] 33.27 29.44 0.832 0.714 0.012 0.030 32.03 25.44 0.902 0.744 0.010 0.046
ReconFormer [14] 33.75 30.42 0.837 0.728 0.011 0.026 32.46 26.47 0.906 0.766 0.010 0.039
FpsFormer [35] 33.74 30.63 0.841 0.732 0.011 0.026 32.35 26.65 0.897 0.768 0.010 0.038
LMO [33] 34.49 31.10 0.846 0.744 0.011 0.023 35.35 27.99 0.921 0.787 0.006 0.028
@@
IV Experiments
IV-A Experimental Settings
IV-A1 Datasets and Evaluation Metrics.
We evaluate HiFi-Mamba on two public MRI datasets: fastMRI (knee) [6] and CC359 (brain) [36]. The fastMRI dataset comprises 1,172 complex-valued single-coil coronal knee scans acquired with Proton Density (PD) weighting. Each scan contains approximately 35 coronal slices with a matrix size of 320×320. We exclusively use the Proton Density Fat-Suppressed (PDFS) subset for both training and evaluation, following the official data split. The CC359 dataset contains raw brain MR scans acquired using clinical MR scanners (Discovery MR750; GE Healthcare, USA). Following the official split, 25 subjects are used for training and 10 for testing. Each slice has a matrix size of 256×256. To eliminate slices with limited anatomical content, we discard the first and last five slices for fastMRI and the first and last 15 slices for CC359.
In our experiments, the inputs are generated by applying equispaced 1D Cartesian undersampling masks, as provided by the fastMRI challenge [6]. Specifically, for an acceleration factor (AF) of 4, the central 8% of k-space lines are fully sampled; for AF=8, this portion is reduced to 4%.
We evaluate reconstruction performance using three widely used metrics: PSNR (Peak Signal-to-Noise Ratio) [37], SSIM (Structural Similarity Index) [38], and NMSE (Normalized Mean Squared Error) [39].
IV-A2 Data Preprocessing Strategy.
@@ IV-A Experimental Settings
IV-B Comparison with State-of-the-Art Methods
We compare HiFi-Mamba against representative state-of-the-art methods spanning three major model paradigms: CNN-based (UNet, ISTA-Net), Transformer-based (ReconFormer, FPSFormer), and Mamba-based (LMO). Evaluations are performed using equispaced 1D Cartesian undersampling at acceleration factors (AF) of 4× and 8× on the fastMRI and CC359 datasets. The quantitative results are summarized in Table I.
HiFi-Mamba (P2), where P2 denotes a patch size of 2, consistently outperforms existing methods across most evaluation scenarios. Notably, it achieves significant gains at AF=8 on both datasets, e.g., reaching 31.38 dB PSNR and 0.758 SSIM on fastMRI, and 28.08 dB PSNR and 0.802 SSIM on CC359. Although its PSNR (34.47 dB) on fastMRI at AF=4 is slightly lower than LMO (34.49 dB), HiFi-Mamba (P2) still surpasses LMO in SSIM and NMSE and consistently outperforms all other methods across the remaining settings, demonstrating strong generalization capability.
@@ IV-B Comparison with State-of-the-Art Methods
The finer-grained variant HiFi-Mamba (P1), with a patch size of 1, further improves reconstruction performance, establishing a new state-of-the-art. Specifically, it achieves 34.85 / 31.81 dB PSNR and 0.855 / 0.762 SSIM on fastMRI (AF=4 / 8), with corresponding NMSE of 0.010 / 0.020. On CC359, it achieves 36.93 dB PSNR and 0.942 SSIM at AF=4, and maintains robust performance at AF=8, reaching 28.49 dB PSNR, 0.810 SSIM, and 0.026 NMSE, consistently surpassing prior state-of-the-art methods across all metrics.
@@
IV-B1 Visualization Results.
Figure 3 presents qualitative comparisons under 4× and 8× acceleration on representative slices from the fastMRI (knee) and CC359 (brain) datasets. The top row shows the reconstructed images, while the bottom row displays the corresponding error maps (difference from ground truth), color-coded from 0 (blue) to 0.2 (red). Yellow circles and red arrows indicate discrepancies in structural detail, while blue boxes mark zoomed-in regions for closer inspection.
ISTA-Net and UNet exhibit edge blurring and loss of structural detail. ReconFormer and FPS-Former partially alleviate these issues but still fail to reconstruct fine anatomical features with high acceleration factors. Notably, FPS-Former produces visually sharp contours in $AF=8$ brain reconstructions, but introduces hallucinated structures inconsistent with the ground truth. The Mamba-based LMO also exhibits boundary degradation and missing details.
@@
Overall, these visual results highlight the effectiveness of our frequency-aware dual-stream design in jointly modeling global context and localized high-frequency information, enabling perceptually accurate and structurally robust reconstructions under high undersampling conditions.
IV-B2 Efficiency
To evaluate the computational efficiency, we report FLOPs and reconstruction accuracy on the fastMRI dataset (AF=8) with $\times320$ resolution images, measured on a single NVIDIA A100 GPU. As shown in Table II, HiFi-Mamba achieves a favorable balance between performance and efficiency. Notably, HiFi-Mamba(P2) delivers strong reconstruction quality (31.38 dB PSNR, 0.758 SSIM) while requiring only 67.87G FLOPs—substantially lower than all competing models. Despite operating with single directional scanning, it matches or surpasses the accuracy of computationally heavier models such as ReconFormer and FpsFormer, both of which rely on computationally intensive attention mechanisms. HiFi-Mamba(P1), although computationally heavier (270.37G FLOPs), achieves the best overall performance (31.81 dB PSNR, 0.762 SSIM), significantly surpassing the state-of-the-art LMO (484.98G FLOPs) in both accuracy and efficiency. These results highlight the scalability and resource-awareness of our architecture, enabling high-fidelity MRI reconstruction with superior performance at reduced computational cost.
IV-C Ablation Studies and Analysis
TABLE III: Model component experiment is conducted on the CC359 dataset with AF =8=8, and patch size =2=2.
$\mathcal{W}$-Lap. HiFi-Mamba DFSA CRM PSNR SSIM NMSE
@@ IV-C Ablation Studies and Analysis
IV-C1 Ablation on Mamba Unit.
We perform an ablation study on the CC359 dataset with AF=8=8 to assess the contribution of each component. Starting with the proposed $\mathcal{W}$-Laplacian transform, we obtain 26.22 dB PSNR and 0.781 SSIM. Incorporating the HiFi-Mamba block, which enables frequency-aware interaction, increases the PSNR to 27.07 dB and the SSIM to 0.790. Adding the DFSA module refines the fused features, improving performance to 27.99 dB PSNR and 0.799 SSIM. Finally, appending the CRM module enhances high-frequency stream representations and achieves the best performance: 28.07 dB PSNR, 0.802 SSIM, and 0.027 NMSE. These results demonstrate the complementary benefits of each module and their collective contribution to reconstruction quality.
TABLE IV: Model architecture experiment is conducted on the CC359 dataset with AF =8=8.
@@
2 3x2 27.92 0.800 0.028
2 4x2 27.95 0.800 0.028
2 6x2 28.07 0.802 0.027
2 8x2 28.15 0.805 0.027
4 6x2 27.71 0.793 0.029
1 6x2 28.49 0.810 0.026
@@
We perform an ablation study on model depth and patch size under an $8\times$ acceleration setting using the CC359 dataset. As shown in Table V, with patch size fixed at 2, increasing depth from $3\times 2$ to $6\times 2$ consistently improves PSNR and SSIM, indicating better anatomical modeling. Further increasing to $8\times 2$ offers only marginal gains, suggesting saturation beyond moderate depth.
@@
We evaluate the impact of kernel size in the spatially-aware refinement module by varying the receptive field of the depthwise 1D convolution applied to the conditioned parameters ($B$, $C$, $\Delta$). As shown in Table V, increasing the kernel size from $3$ to $7$ yields consistent improvements across all metrics. The $7\times 7$ configuration achieves the best performance, indicating that larger spatial context enhances the anatomical coherence of the learned dynamics.
TABLE V: Spatially-Aware experiment is conducted on the CC359 dataset with AF =8=8, and patch size =1=1.
@@
DConv1D ($3\times 3$) 27.81 0.796 0.030
DConv1D ($5\times 5$) 28.05 0.805 0.028
DConv1D ($7\times 7$) 28.49 0.810 0.026
V Conclusion

### v2

@@ Abstract
Codehttps://github.com/HongliChen-Ali/HiFi-Mamba.git
@@
Introduction
Magnetic Resonance Imaging (MRI) is a clinically indispensable modality due to its non-invasive nature and excellent soft-tissue contrast (Jerban et al. 2020; Varela-Mattatall and D’Arcy 2017). However, a primary limitation lies in its long acquisition time, which can cause patient discomfort and increase the risk of motion artifacts (Hammernik et al. 2018; Knoll et al. 2020). To accelerate scans, modern protocols commonly adopt undersampling in the frequency domain. While this reduces scan time, it violates the Nyquist-Shannon sampling criterion (Ye 2019; Zbontar et al. 2018), leading to aliasing artifacts and degraded image quality. Addressing this challenge requires advanced reconstruction methods capable of recovering high-fidelity images from incomplete k-space data. Traditional compressed sensing (CS) methods (Lustig et al. 2007) exploit sparsity priors in transform domains (e.g., wavelets), but they require extensive hyperparameter tuning and often lack robustness to variations in sampling patterns or noise conditions.
@@ Introduction
Recent deep learning frameworks have substantially advanced MRI reconstruction by leveraging data-driven priors from large-scale datasets (Zhou et al. 2018). Convolutional neural networks (CNNs) have demonstrated strong performance by modeling hierarchical and localized anatomical features (Schlemper et al. 2018; Qin et al. 2019; Chen et al. 2024). Model-based CNNs further improve reconstruction quality by integrating the MRI forward model and enforcing data consistency (Hammernik et al. 2018; Aggarwal et al. 2019; Shan et al. 2024). However, the inherent locality of CNNs limits their capacity to capture long-range dependencies, which are crucial for reconstructing global anatomical structures, especially under highly undersampled conditions.
@@
Transformer-based architectures model global dependencies through self-attention by computing pairwise correlations across all spatial tokens (Vaswani et al. 2017; Fu et al. 2025). This capacity has shown promise in MRI reconstruction (Guo et al. 2023; Wu et al. 2023) by enabling global context modeling to restore anatomical structures. Nonetheless, standard Transformers incur quadratic complexity, posing challenges for high-resolution MRI applications. To improve efficiency, variants such as the Swin Transformer (Liu et al. 2021) employ hierarchical designs and shifted window-based attention to restrict computations to local neighborhoods. While computationally efficient, these approaches inherently constrain the receptive field, reigniting the fundamental trade-off between computational efficiency and global modeling capability, leaving a critical gap for a more holistic solution.
@@
Structured State Space Models (SSMs) (Gu et al. 2022) have recently emerged as promising alternatives to self-attention for global context modeling, offering linear-time scalability with strong sequence modeling capacity. Among them, Mamba (Gu et al. 2023) employs input-dependent state transitions and efficient gating mechanisms, enabling expressive long-range interactions with reduced computational complexity. While initially proposed for language tasks, Mamba is rapidly emerging as a powerful alternative in visual domains such as image classification and restoration (Guo et al. 2024; Liu et al. 2024).
However, a direct application of Mamba to MRI reconstruction reveals three fundamental limitations rooted in its original design. First, existing vision-specific Mamba architectures utilize multi-directional spatial scanning (Rahman et al. 2024; Zhang et al. 2024) to enhance coverage, but this introduces significant computational redundancy (see Figure 1(a)).
@@
Second, its state-space parameters are derived independently for each spatial token through local transformations, limiting spatial awareness. This design neglects neighboring context, which is essential for modeling coherent anatomical structures in MRI. Third, existing Mamba-based variants lack sensitivity to high-frequency components (Ma et al. 2024), which are critical for preserving fine anatomical details in MRI reconstruction.
@@
CNNs have been widely employed in MRI reconstruction for their ability to extract hierarchical features efficiently. Early models such as DeepCascade (Schlemper et al. 2018) and ISTA-Net (Zhang and Ghanem 2018) framed reconstruction as an unrolled optimization process, integrating data fidelity with learnable, network-based priors. Subsequent approaches, including KIKI-Net (Eo et al. 2018) and DuDoRNet (Zhou et al. 2020), incorporated architectural advances such as residual connections, dilated convolutions, and hybrid modeling across image and k-space domains to enhance reconstruction quality and optimization stability. However, the inherently localized receptive fields of CNNs constrain their ability to capture global anatomical context. This limitation has motivated the development of architectures with enhanced global modeling capacity.
@@
Transformers have been increasingly adopted in MRI reconstruction for their ability to model long-range dependencies through global self-attention, which facilitates the preservation of anatomical structures. Early models such as SLATER (Korkmaz et al. 2022) and TTM (Gu et al. 2021) applied Transformer blocks in the image domain. DuDReTLU-Net (Wang et al. 2023) jointly models image and $k$-space domains through a transformer-based architecture to enhance reconstruction quality. Later variants such as SwinMR (Huang et al. 2022) and ReconFormer (Guo et al. 2023) adopted hierarchical and windowed self-attention mechanisms to improve multi-scale feature modeling and computational efficiency. Although these designs reduce computational overhead, they often rely on localized attention and staged aggregation, which may still limit the capacity for global context modeling in high-resolution MRI reconstruction.
Structured State Space Models and Mamba.
State space models (SSMs) have recently gained attention for their ability to model long-range dependencies with linear complexity (Li et al. 2024). The vision-centric Mamba variant, vMamba (Liu et al. 2024), introduced a four-directional scanning strategy to enhance 2D spatial context modeling. Recent works, such as MambaMIR (Huang et al. 2025) and LMO (Li et al. 2025), have extended this paradigm for MRI reconstruction. However, these adaptations remain suboptimal in both efficiency and fidelity. Their reliance on multi-directional scanning induces considerable computational overhead due to repeated processing of spatial features—an issue exacerbated in high-resolution scenarios. Moreover, they inherit two core limitations of the original Mamba design: purely local state-space parameterization and insensitivity to high-frequency anatomical details. These constraints highlight a gap between current vision-based Mamba variants and the spectral characteristics of MRI.
Overall Pipeline
Our proposed HiFi-Mamba network follows an unrolled optimization framework, a powerful paradigm for solving inverse problems like MRI reconstruction. The network backbone consists of $K=8$ cascaded HiFi-Mamba Groups. As shown in Figure 2, each Group consists of two sequential Mamba Units for feature refinement, followed by a Data Consistency (DC) block. Specifically, the input undersampled image $I_{\text{in}}\in\mathbb{R}^{H\times W\times 2}$ is a two-channel tensor representing the real and imaginary parts of the complex-valued MR data. The input tensor is first transformed into patch embeddings as:
@@
$\mathcal{w}$-Laplacian Block
@@
While the Laplacian pyramid (Burt and Adelson 1987) offers residual-based multiscale representations, its decomposition is resolution-oriented rather than frequency-structured. The resulting low-frequency component is a blurred, downsampled approximation without explicit spectral semantics, limiting its utility for preserving anatomical context in MRI. To overcome this limitation, we replace the resolution-based hierarchy with a wavelet-based formulation that enables structured and reversible frequency separation.
@@
HiFi-Mamba Block
@@
[…]
$times 4×\times 8$$\times$ $times 4×\times 8$$\times$ $times 4×\times 8$$\times$
@@
UNet (Zhou et al. 2018) 31.66 28.60 0.798 0.697 0.021 0.035 28.27 24.28 0.847 0.720 0.025 0.059
ISTA (Zhang and Ghanem 2018) 33.27 29.44 0.832 0.714 0.012 0.030 32.03 25.44 0.902 0.744 0.010 0.046
ReconFormer (Guo et al. 2023) 33.75 30.42 0.837 0.728 0.011 0.026 32.46 26.47 0.906 0.766 0.010 0.039
FpsFormer (Meng et al. 2025) 33.74 30.63 0.841 0.732 0.011 0.026 32.35 26.65 0.897 0.768 0.010 0.038
LMO (Li et al. 2025) 34.49 31.10 0.846 0.744 0.011 0.023 35.35 27.99 0.921 0.787 0.006 0.028
@@
Table 1: Quantitative results on fastMRI-Equispaced and CC359-Equispaced under $4\times$ and $8\times$ accelerations.
Experiments
Experimental Settings
Datasets and Evaluation Metrics.
We evaluate HiFi-Mamba on two public MRI datasets: fastMRI (knee) (Zbontar et al. 2018) and CC359 (brain) (Warfield et al. 2004). The fastMRI dataset comprises 1,172 complex-valued single-coil coronal knee scans acquired with Proton Density (PD) weighting. Each scan contains approximately 35 coronal slices with a matrix size of $320\times 320$. We exclusively use the Proton Density Fat-Suppressed (PDFS) subset for both training and evaluation, following the official data split. The CC359 dataset contains raw brain MR scans acquired using clinical MR scanners (Discovery MR750; GE Healthcare, USA). Following the official split, 25 subjects are used for training and 10 for testing. Each slice has a matrix size of $256\times 256$. To eliminate slices with limited anatomical content, we discard the first and last five slices for fastMRI and the first and last 15 slices for CC359.
In our experiments, the inputs are generated by applying equispaced 1D Cartesian undersampling masks, as provided by the fastMRI challenge (Zbontar et al. 2018). Specifically, for an acceleration factor (AF) of 4, the central 8% of k-space lines are fully sampled; for $\text{AF}=8$, this portion is reduced to 4%.
We evaluate reconstruction performance using three widely used metrics: PSNR (Peak Signal-to-Noise Ratio) (Huynh-Thu and Ghanbari 2008), SSIM (Structural Similarity Index) (Wang et al. 2004), and NMSE (Normalized Mean Squared Error) (Zhao et al. 2016).
Data Preprocessing Strategy.
@@
Comparison with State-of-the-Art Methods
We compare HiFi-Mamba against representative state-of-the-art methods spanning three major model paradigms: CNN-based (UNet, ISTA-Net), Transformer-based (ReconFormer, FPSFormer), and Mamba-based (LMO). Evaluations are performed using equispaced 1D Cartesian undersampling at acceleration factors (AF) of 4× and 8× on the fastMRI and CC359 datasets. The quantitative results are summarized in Table 1.
HiFi-Mamba (P2), where P2 denotes a patch size of 2, consistently outperforms existing methods across most evaluation scenarios. Notably, it achieves significant gains at $\text{AF}=8$ on both datasets, e.g., reaching 31.38 dB PSNR and 0.758 SSIM on fastMRI, and 28.08 dB PSNR and 0.802 SSIM on CC359. Although its PSNR (34.47 dB) on fastMRI at $\text{AF}=4$ is slightly lower than LMO (34.49 dB), HiFi-Mamba (P2) still surpasses LMO in SSIM and NMSE and consistently outperforms all other methods across the remaining settings, demonstrating strong generalization capability.
@@
Table 2: Efficiency comparison on fastMRI ($\text{AF}=8$, $320\times 320$) using NVIDIA A100.
The finer-grained variant HiFi-Mamba (P1), with a patch size of 1, further improves reconstruction performance, establishing a new state-of-the-art. Specifically, it achieves $34.85$ / $31.81$ dB PSNR and $0.855$ / $0.762$ SSIM on fastMRI ($\text{AF}=4$ / $8$), with corresponding NMSE of $0.010$ / $0.020$. On CC359, it achieves 36.93 dB PSNR and 0.942 SSIM at $\text{AF}=4$, and maintains robust performance at $\text{AF}=8$, reaching 28.49 dB PSNR, 0.810 SSIM, and 0.026 NMSE, consistently surpassing prior state-of-the-art methods across all metrics.
@@
Figure 3: Qualitative comparison on the fastMRI and CC359 datasets under single-coil settings. (a) Reconstruction results on the fastMRI knee dataset with acceleration factors $\text{AF}=4$ and $\text{AF}=8$. (b) Reconstruction results on the CC359 brain dataset under the same acceleration factors. The second row of each subplot shows the corresponding error maps. The blue boxes, yellow ellipses and red arrow highlight the details in the reconstruction results.
Visualization Results.
Figure 3 presents qualitative comparisons under $4\times$ and $8\times$ acceleration on representative slices from the fastMRI (knee) and CC359 (brain) datasets. The top row shows the reconstructed images, while the bottom row displays the corresponding error maps (difference from ground truth), color-coded from 0 (blue) to 0.2 (red). Yellow circles and red arrows indicate discrepancies in structural detail, while blue boxes mark zoomed-in regions for closer inspection.
ISTA-Net and UNet exhibit edge blurring and loss of structural detail. ReconFormer and FPS-Former partially alleviate these issues but still fail to reconstruct fine anatomical features with high acceleration factors. Notably, FPS-Former produces visually sharp contours in $\text{AF}=8$ brain reconstructions, but introduces hallucinated structures inconsistent with the ground truth. The Mamba-based LMO also exhibits boundary degradation and missing details.
@@
Efficiency
We evaluate computational efficiency on the fastMRI dataset ($\text{AF}=8$, $320\times 320$) using a single NVIDIA A100 GPU. As shown in Table 2, HiFi-Mamba achieves a favorable balance between performance and efficiency. HiFi-Mamba (P2) achieves 31.38 dB PSNR and 0.758 SSIM with only 67.87G FLOPs—substantially lower than all competing models, while matching or surpassing heavier attention-based architectures such as ReconFormer and FpsFormer. HiFi-Mamba (P1), though requiring 270.37G FLOPs, delivers the best overall performance (31.81 dB PSNR, 0.762 SSIM), outperforming the state-of-the-art LMO (484.98G FLOPs) in both accuracy and efficiency. These results demonstrate the resource-awareness of our design for high-fidelity MRI reconstruction.
Ablation Studies and Analysis
[…]
@@
Table 3: Ablations on CC359 ($\text{AF}=8$, $\text{patch}=2$).
Ablation on Mamba Unit.
We perform an ablation study on the CC359 dataset with $\text{AF}=8$ to assess the contribution of each component. Starting with the proposed $\mathcal{w}$-Laplacian transform, we obtain 27.07 dB PSNR and 0.790 SSIM. Incorporating the HiFi-Mamba block, which enables frequency-aware interaction, increases the PSNR to 27.46 dB and the SSIM to $0.794$. Adding the DFSA module refines the fused features, improving performance to 27.99 dB PSNR and 0.799 SSIM. Finally, appending the CRM module enhances high-frequency stream representations and achieves the best performance: 28.07 dB PSNR, 0.802 SSIM, and 0.027 NMSE. These results demonstrate the complementary benefits of each module and their collective contribution to reconstruction quality.
@@
2 $3\times 2$ 27.92 0.800 0.028
2 $4\times 2$ 27.95 0.800 0.028
2 $6\times 2$ 28.07 0.802 0.027
2 $8\times 2$ 28.15 0.805 0.027
4 $6\times 2$ 27.71 0.793 0.029
1 $6\times 2$ 28.49 0.810 0.026
Table 4: Model architecture experiment is conducted on the CC359 dataset with $\text{AF}=8$.
@@
We perform an ablation study on model depth and patch size under an $8\times$ acceleration setting using the CC359 dataset. As shown in Table 4, with patch size fixed at 2, increasing depth from $3\times 2$ to $6\times 2$ consistently improves PSNR and SSIM, indicating better anatomical modeling. Further increasing to $8\times 2$ offers only marginal gains, suggesting saturation beyond moderate depth.
@@
We evaluate the impact of kernel size in the spatially-aware refinement module by varying the receptive field of the depthwise 1D convolution applied to the conditioned parameters ($B$, $C$, $\Delta$). As shown in Table 5, increasing the kernel size from $3$ to $7$ yields consistent improvements across all metrics. The kernel size 7 configuration achieves the best performance, indicating that larger spatial context enhances the anatomical coherence of the learned dynamics.
@@
DConv1D ($k=3$) 27.81 0.796 0.030
DConv1D ($k=5$) 28.05 0.805 0.028
DConv1D ($k=7$) 28.49 0.810 0.026
Table 5: Spatially-Aware experiment is conducted on the CC359 dataset with $\text{AF}=8$, and $\text{patch size}=1$. Here, $k$ denotes the kernel size.
Conclusion